A body is thrown vertically upwards with initial velocity of 100m/s.Find 1.time taken to reach at maximum height 2.maximum height attend 3.velocity after 5 sec 4.velocity after 16 sec. Take a=10m/s2.
1
Expert's answer
2012-08-09T10:48:48-0400
1. Here is the formula for the height of a body:

H(t) = Vt - gt²/2

Let's derivate it by t:

H'(t) = V - gt

So,

V - gt = 0 ==> t = V/g = 100 [m/s] / 10 [m/s²] = 10 s;

2. H(10) = 100*10 - 10*10²/2 = 500;

3. V(5) = H'(5) = V - 5g = 100 - 50 = 50 m/s;

4. V(16) = H'(5) = V - 16g = 100 - 160 = -60 m/s.